OpenStudy (anonymous):

Pleaseeeee help... x/4 + 2x/3=11..

OpenStudy (jhannybean):

\[\large \frac{x}{4}+\frac{2x}{3}= 11\] first we need a common denominator,so we multiply both fractions by 4 and 3.

OpenStudy (anonymous):

yeah i got 24 for the common denominator.

OpenStudy (jhannybean):

\[\large \frac{3}{3}*\frac{x}{4}+ \frac{4}{4}{*\frac{2x}{3}}=11\]\[\large \frac{3x}{12}+\frac{8x}{12} = 11\]now that our denominators are the same,wecan rewrite it as one fraction \[\large \frac{3x+8x}{12} = 11\]multiply by the reciprocal which is (1/12) or cross multiply to get rid of the fraction on the LHS of the equation \[\large 3x+8x= 132\] combine the stuff on the LHS of the equation \[\large 11x = 132\] divide both sides by 11 \[\large \frac{\cancel {11}x}{\cancel{11}}= \frac{132}{11}\] simplify and you'll have your answer :)

OpenStudy (jhannybean):

sorry, multiplying by the reciprocal would mean multiplying by (12) not (1/12) *** sorry about that.

OpenStudy (anonymous):

oh ok. thanks

OpenStudy (jhannybean):

Do you understand how to solve for x now?

OpenStudy (anonymous):

yeah, i do. thank youuu. you wanna give me a practice problem so i can prove myself. lol

OpenStudy (jhannybean):

hmm...pull out another problem from your book and i'll help you work through it :P

OpenStudy (anonymous):

okie i got one. x/3 + 3x/4 = 13

OpenStudy (anonymous):

i got x=39

OpenStudy (jhannybean):

hmmm...let me solve it :P did you use the above steps? ^

OpenStudy (anonymous):

yes mam

OpenStudy (jhannybean):

i'm not getting 39.... :\ what did you get as your denominator?

OpenStudy (anonymous):

i used 12

OpenStudy (jhannybean):

and for your numerator?

OpenStudy (anonymous):

i left them alone.

OpenStudy (anonymous):

here lemme draw out what i did.

OpenStudy (jhannybean):

lol wait aminute. maybe i did it wrong -_- retrying! :P

OpenStudy (anonymous):

oh okay

OpenStudy (jhannybean):

x/3+ 3x/4 = 13 4x/ 12 + 9x/12 =13 (4x+9x)= (12)(13) 13x = 156 x= ?

OpenStudy (anonymous):

oh i didnt know we were suppose to do anything with the numerators. since we didnt in the previous problem

OpenStudy (jhannybean):

correction: look at the 4rth step :P

OpenStudy (anonymous):

OHH. I GET IT NOW. haha my bad. thanks

OpenStudy (jhannybean):

xD

OpenStudy (anonymous):

i recommend you stay online cuz im probably gonna need some more help. :P

OpenStudy (anonymous):

and your like my only hope to passing

OpenStudy (jhannybean):

lol, are the types of problems you're having trouble with like these?

OpenStudy (jhannybean):

you can always "fan" be by left clicking on my name and "becoming a fan" that way you can message me with your problems! ^_^

OpenStudy (jhannybean):

me*

OpenStudy (anonymous):

Good idea! done. There mostly inequalities, but theres some old geometry stuff that i would like to go over to make sure i can do. Like sine, cosine, tangent. That kinda stuff :P

OpenStudy (jhannybean):

ohh okay. just post up a new question after closing this one, or send me a message with it xD I'll try my best to help!

OpenStudy (anonymous):

okie thanks! your awesomee haha

OpenStudy (jhannybean):

and if you liked my responses there is a "best response" button to the right :) --->

OpenStudy (anonymous):

oh okay! haha. already clicked. wooo

OpenStudy (jhannybean):

:D